Grade 12th passMechanics

A stone dropped from the top of a tower is found to travel 5/9 of the height of the tower during the last second of fall.....What is the total time of fall?

Saikat
8 Years agoGrade 12th pass

1 Answer

abhay bajaj
8 Years ago
Let the height of tower be xu = 0m/sso let time to cover 5/9x of building be t secso time taken to cover 4x/9 of the building = t - 1 sec **S_{t} = u + 1/2g(2t - 1)** 5x/9= 5(2t - 1) ...........(i) g = 10so again s = ut +1/2gt²4x/9 = 0 + 5(t - 1)² ..................(ii) dividing i and ii 5x/9/4x/9 =5(2t - 1) /5(t - 1)²⇒5t² +5 - 10t = 8t - 4⇒5t² -18t +9 =0solving t we get t = 3 sec and 0.6 sec